Direct Materials Mix and Yield Variances Chypre, Inc., produces a cologne mist using a solvent mix (water and pure alcohol) aRequired: Instructions for parts 3 and 4: If a variance is zero, enter 0 and select Not applicable from the drop down boxPlease help question No. 3 thanks!
Answer #1

Material Yield Variance = (Standard Qty - Actual Qty)*Standard cost per unit

Standard Qty = 900*0.95 = 855 gallons

Material Yield Variance = (855-775)*435.776596

Material Yield Variance = 34862.1